OpenStudy (anonymous):

Heidi and Mike have 51 dimes and nickles. If the value of the coins is $4.10, how many coins of each type were there?

OpenStudy (lgbasallote):

let x = no. of nickles y = no. of dimes x + y = 51 5(cents)x + 10(cents)y = 410 cents multiply the first equation by -10.. -10x - 10y = -510 5x + 10y = 410 -5x = -100 x = 20 y = 31 There are 20 nickles and 31 dimes

OpenStudy (anonymous):

thank you for being so specific! :)